# HoPE: Hybrid of Position Embedding for Length Generalization in Vision-Language Models

### v1

@@ 4.2 Dynamic Temporal Scaling Mechanism
Considering the distinct information densities of text and visual tokens, HoPE introduces a dynamic temporal scaling mechanism that adjusts the temporal strides of visual inputs. Specifically, we first define a set of scaling factors, e.g., $\Gamma=\{0.5,0.75,1,1.25,1.5\}$, which includes both stretching ($\gamma>1$) and compressing $(\gamma<1$) operations. During training, scaling factor $\gamma$ is randomly selected from $\Gamma$ and applied to each video. This allows the model to learn temporal relationships at multiple scales, making it more robust to variations in video speed, which are common in real-world scenarios. Consider a multimodal input $(text,video,text)$ of length $L_{t}$, $L_{v}$, and $L_{e}$, respectively. The position indices $(t,x,y)$ for each token with our dynamic scaling factor ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}$ are:
$(t,x,y)=\begin{cases}(l,l,l),&0\leq l<L_{t}\\ \\ \begin{pmatrix}L_{t}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}(l-L_{t}),\\ L_{t}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}(l-L_{t})+w-\frac{W}{2},\\ L_{t}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}(l-L_{t})+h-\frac{H}{2}\end{pmatrix},&L_{t}\leq l<L_{t}+L_{v}\\ \\ \begin{pmatrix}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}-1)L_{v}+l,\\ ({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}-1)L_{v}+l,\\ ({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}-1)L_{v}+l\end{pmatrix},&L_{t}+L_{v}\leq l<L_{t}+L_{v}+L_{e}\end{cases}.$ (7)
During inference, scaling factors can be flexibly selected from the set to accommodate videos of different lengths. It is worth noting that unlike existing methods, which do not consider temporal scaling for visual tokens [1, 2, 4, 5] or just apply a fixed and unidirectional scaling factor for both training and testing [23], our methods not only help the model learn temporal relationships at multiple scales, but also offer flexibility during inference to accommodate various context lengths.
@@ 5 Experiment
In this section, we evaluate the performance of HoPE on four video benchmarks across long video understanding and long video retrieval tasks, aiming to validate its effectiveness in multimodal long-context modeling. Additionally, we conduct ablation studies to investigate the individual contribution of each strategy to overall performance.
@@ 5.1 Experimental Setups
Implementation Details. We utilize Qwen2-1.5B and Qwen2-7B [33] as the backbone models. By integrating these models with vision encoders from Qwen2-VL-2B/7B-Instruct [2], we obtain Qwen2-2B/7B-Video, respectively. During training, we adopt a batch size of 128, a learning rate of 1e-5(2B)/2e-5(7B) with a cosine scheduler. Following the instruction tuning settings in Qwen2-VL [2], we set the maximum video frames to 128 and the video sampling rate to 2. The training context length is set to 8k, with the entire training process taking approximately 304 GPU hours on machines equipped with H800-80GB GPUs. During evaluation, the minimum tokens per frame is set to 144.
Training Data. We train the models on a subset of LLaVA-Video-178k [34], which consists of 178k videos ranging from 0 to 3 minutes and 5M instruction samples, including captions, free-form and multiple-choice question answering. Our selected subset includes 30k videos with durations under 2 minutes and 3k videos with durations between 2 and 3 minutes, resulting in roughly 300k pairs.
Baselines. We compare HoPE with the following RoPE-variants: 1) vanilla RoPE [18], the standard approach in long-context LLMs, 2) M-RoPE [2], a famous RoPE extension in Qwen2-VL for multimodal inputs, 3) VideoRoPE [23], a specialized RoPE variant designed for video-related tasks.
Evaluation Benchmarks. We evaluate HoPE across four video benchmarks for long video understanding and long video retrieval tasks. For long video understanding, we utilize LongVideoBench [35], Video-MME [36], and MLVU [37]. For long video retrieval, we employ V-NIAH (Visual Needle-In-A-Haystack) [12]. For detailed benchmark descriptions, please refer to Appendix B.1.
Table 1: Performance comparison on long video understanding task across MLVU, LongVideoBench, and Video-MME benchmarks. The training context length for all methods is set to 8k, while we report the performance on 8k, 16k, 32k, and 64k to assess the length generalization ability of different methods. The best results are bold, while the second best results are underlined.
@@ 5.2 Results on Long Video Understanding
In this section, we provide a comprehensive comparison of HoPE and different RoPE variants in long video understanding. From Table 1, we observe that: (1) HoPE consistently outperforms all baselines across nearly all benchmarks, context lengths, and backbone sizes. Specifically, under the 7B model scale and 32k context lengths, HoPE surpasses vanilla RoPE by 2.82, 4.05, 1.45 on MLVU, LongVideoBench, and Video-MME, respectively. This confirms its effectiveness and generalizability in multimodal long-context modeling. (2) The effectiveness of HoPE scales with backbone size. For instance, when the size of the backbone LLM increases from 2B to 7B, HoPE’s performance gain on LongVideoBench (32k) significantly increases from 0.66 to 4.05 compared to vanilla RoPE. Notably, the performance gap between different methods on the 2B scale is less significant, probably due to the limited capabilities of the backbone LLM. (3) For context lengths under 64k, performance on Video-MME drops substantially, while the impact on MLVU and LongVideoBench is less pronounced. This suggests that extrapolating to extreme context lengths (e.g., up to 8x) remains highly challenging.
Figure 3: Performance on long video retrieval task (V-NAIH). The black dotted line represents the training context length. Each frame corresponds to 144 tokens.
@@ 5.3 Results on Long Video Retrieval
5.4 Ablation Studies
Table 2: Ablation results on Video-MME from 8k to 64k. Here, HFA: hybrid frequency allocation, DTS: dynamic temporal scaling.
@@ 5.4 Ablation Studies
+ 3D + HFA 57.66 59.19 59.33 26.98
+ 3D + HFA + DTS 57.74 59.33 59.44 27.34
We conduct a series of ablation experiments to evaluate the impact of each component in HoPE and summarize the results in Table 2. According to the results, we observe that : (1) The 3D structure effectively improves the performance of vanilla RoPE in multimodal contexts, supporting our Proposition 3.1. (2) Based on the 3D structure, the hybrid frequency allocation (HFA) strategy further enhances long-range semantic modeling, achieving an average improvement of 1.69 across all context lengths. (3) The dynamic temporal scaling (DTS) mechanism facilitates VLMs’ robustness to varying video speeds in real-world scenarios, yielding further performance gain. By combining the above strategies, our HoPE achieves the best results in multimodal long-context modeling. More ablations on test-time scaling factor selection are provided in Appendix B.3.
@@ 6 Related Work
Position Embeddings in LLMs. Rotary Position Embedding (RoPE) has become a common choice in LLMs [19, 20, 21]. As discussed in Section 2, RoPE achieves this success through rotating query and key vectors, encoding relative position information through an absolute positional encoding approach. Despite its success, several works have pointed out that No Position Embedding (NoPE) still works for decoder-only LLMs, arguing that the causal attention mechanism implicitly learns absolute position information [25, 26, 38]. These works even suggest NoPE outperforms RoPE in out-of-distribution (OOD) scenarios. However, this observation remains unexplored in multimodal settings, where positional encoding strategies may have different implications for cross-modal interactions.
Multimodal Position Embeddings in VLMs. In current VLMs [1, 2, 3, 4, 5], images are processed by vision encoders and flattened into 1D tokens. Some of these VLMs [1, 4, 5] still use vanilla RoPE to encode position information, which distorts spatial-temporal locality. Recently, Qwen2-VL [2] has proposed M-RoPE, which extends RoPE by assigning different frequencies to encode different positional components. VideoRoPE [23] further refines this method by adjusting frequency allocation strategies and applying a fixed scaling factor for the temporal indices of visual tokens. In contrast, our work theoretically analyzes how different frequency allocation strategies impact the performance of multimodal RoPE. Based on our analysis, we propose HoPE, which combines RoPE and NoPE in the multimodal scenario for length generalization and demonstrates strong empirical performance.
@@ A.2 Semantic Preference Property
We now prove that the frequency allocation strategies in current multimodal RoPEs are unreliable in capturing semantic similarities over extended contexts, as defined in Definition 4.
Definition 4. (Semantic Preference). For any query vector $\mathbf{q}$ and a semantically similar key vector $\mathbf{k}^{\prime}$ that can be expressed as $\mathbf{k}^{\prime}=\mathbf{q}+\delta$ where $\delta$ is a zero-mean perturbation, the attention score with RoPE should satisfy:
@@
2. Monotonicity once the first zero enters. As soon as $\theta\geq\frac{\pi}{2(L-1)}$, the point satisfying $\theta\,\Delta t=\pi/2$ lies in $\{0,\dots,L-1\}$. Each further increase in $\theta$ extends the interval of length $\theta(L-1)$, adding more half-periods of cosine. Each added half-period contains exactly one “negative” region of length $\pi$. Therefore the count |$\bigl{|}\{\Delta:\cos(\theta\Delta)<0\}\bigr{|}$ (and hence $P_{\neg}(\theta)$) can only stay the same or increase, up to $O(1/L)$ rounding errors on the discrete grid.
[…]
@@
Theorem 3.1. Let $X=[x_{1},x_{2},\dots,x_{L}]$ be an input sequence, and let RoPE use any fixed set of temporal frequencies (e.g., highest or lowest). Then there exists a critical length $L_{c}$ such that for all $L\geq L_{c}$, the semantic preference property (Definition 4) is violated.
@@
Assume that each component of the query vector $\mathbf{q}$ is independently and identically distributed with mean $\mu$ and variance $\sigma^{2}$. We denote key vector that is semantically similar to $\mathbf{q}$ as $\mathbf{k}^{\prime}=\mathbf{q}+\delta$, where $\delta$ is a zero-mean perturbation. The semantically unrelated key vector $\mathbf{k}$ is independently drawn with the same distribution as $\mathbf{q}$. Let $\Delta t,\Delta x,\Delta y$ denote relative temporal and spatial distances between the query and each key. According to Definition 4, the semantic preference property requires that:
@@
where $i_{t},i_{x},i_{y}$ denote dimensions allocated to encode temporal $(t)$, horizontal spatial $(x)$, and vertical spatial $(y)$ information. To satisfy the semantic preference property (Definition 4), the expected attention between a query and its semantically similar key should remain higher than that for an unrelated key, regardless of their relative distance. This implies the following condition must hold universally:
@@
However, despite using the lowest frequencies, VideoRoPE still fails to guarantee that the semantic preference property holds for all context lengths (Equation 12). Let VideoRoPE allocate only the smallest frequency to the temporal dimensions, instead of the $n=i_{t}$ smallest frequencies:
@@
$02\sigma^{2}\ |i_{t}|\cos(\Delta t\ \cdot\theta_{\min})<$ 0 .
This single counterexample $(\Delta t,\Delta x,\Delta y)$ violates the semantic preference condition, since no further temporal frequencies are available to “rescue” the sum. Therefore, despite using the lowest frequencies for temporal modeling, VideoRoPE still fails to guarantee the semantic preference property. In conclusion, all frequency allocation strategies in current multimodal RoPEs fail to maintain the semantic preference property in Definition 4, completing the proof. ∎

### v2

@@ 4.2 Dynamic Temporal Scaling Mechanism
Considering the distinct information densities of text and visual tokens, HoPE introduces a dynamic temporal scaling mechanism that adjusts the temporal strides of visual inputs. Specifically, we first define a set of scaling factors, e.g., $\Gamma=\{0.5,0.75,1,1.25,1.5\}$, which includes both stretching ($\gamma>1$) and compressing $(\gamma<1$) operations. During training, the scaling factor $\gamma$ is randomly selected from $\Gamma$ and applied to each video. This allows the model to learn temporal relationships at multiple scales, making it more robust to variations in video speed, which are common in real-world scenarios. Consider a multimodal input $(text,video,text)$ of length $L_{t}$, $L_{v}$, and $L_{e}$, respectively. The position indices $(t,x,y)$ for each token with our dynamic scaling factor $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma$} are:
$\small(t,x,y)=\begin{cases}(l,l,l),&0\leq l<L_{t}\\ \\ \begin{pmatrix}L_{t}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}(l-L_{t}),\\ L_{t}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}(l-L_{t})+w-\frac{W}{2},\\ L_{t}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}(l-L_{t})+h-\frac{H}{2}\end{pmatrix},&L_{t}\leq l<L_{t}+L_{v}\\ \\ \begin{pmatrix}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}-1)L_{v}+l,\\ ({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}-1)L_{v}+l,\\ ({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma}-1)L_{v}+l\end{pmatrix},&L_{t}+L_{v}\leq l<L_{t}+L_{v}+L_{e}\end{cases}.$ (5)
Note that for visual tokens ($L_{t}\leq l<L_{t}+L_{v}$), $l-L_{t}$ indicates the distance of the current frame from the start frame. During inference, scaling factors can be flexibly selected from the set to accommodate videos of different lengths. It is worth noting that unlike existing methods, which do not consider temporal scaling for visual tokens [1, 2, 4, 5] or just apply a fixed and unidirectional scaling factor for both training and testing [28], our methods not only help the model learn temporal relationships at multiple scales, but also offer flexibility during inference to accommodate various context lengths.
@@ 5 Experiment
In this section, we evaluate the performance of HoPE on four video benchmarks across long video understanding and long video retrieval tasks, aiming to validate its effectiveness in multimodal long-context modeling. Additionally, we conduct ablation studies to investigate the individual contribution of each strategy to overall performance and the interplay between task type, context length, and scaling factor selection.
@@ 5.1 Experimental Setups
Implementation Details. We utilize Qwen2-1.5B and Qwen2-7B [39] as the backbone models. By integrating these models with vision encoders from Qwen2-VL-2B/7B-Instruct [2], we obtain Qwen2-2B/7B-Video, respectively. During training, we adopt a batch size of 128, a learning rate of 1e-5(2B)/2e-5(7B) with a cosine scheduler. Following the instruction tuning settings in Qwen2-VL [2], we set the maximum video frames to 128 and the video sampling rate to 2. The training context length is set to 8k, with the entire training process taking approximately 304 GPU hours on machines equipped with H800-80GB GPUs. During evaluation, the minimum tokens per frame are set to 144.
Training Data. We train the models on a subset of LLaVA-Video-178k [40], which consists of 178k videos ranging from 0 to 3 minutes and 5M instruction samples, including captions, free-form, and multiple-choice question answering. Our selected subset includes 30k videos with durations under 2 minutes and 3k videos with durations between 2 and 3 minutes, resulting in roughly 300k pairs.
Baselines. We compare HoPE with the following RoPE variants: 1) vanilla RoPE [23], the standard approach in long-context LLMs, 2) M-RoPE [2], a famous RoPE extension in Qwen2-VL for multimodal inputs, 3) VideoRoPE [28], a specialized RoPE variant designed for video-related tasks.
Evaluation Benchmarks. We evaluate HoPE across four video benchmarks for long video understanding and long video retrieval tasks. For long video understanding, we utilize LongVideoBench [41], Video-MME [42], and MLVU [43], covering videos ranging from a few seconds to 2 hours. For long video retrieval, we employ V-NIAH (Visual Needle-In-A-Haystack) [17]. In this task, a "needle" image is randomly inserted into a "haystack" video, and the VLM is required to answer a question specifically about the embedded "needle" image. Following the protocol in V-NIAH [17], we utilize a haystack video with 1-hour duration (3,000 frames) and insert the needle image at 20% depth intervals (e.g., a frame depth of 0% would place the needle image at the very beginning of the video). For more detailed benchmark descriptions, please refer to Appendix B.1.
Table 1: Performance comparison on long video understanding benchmarks. The training context length of all methods is set to 8k, and we report the performance on 8k, 16k, 32k, and 64k to evaluate length generalization. The best results are bold, while the second best results are underlined.
@@ 5.2 Results on Long Video Understanding
In this section, we provide a comprehensive comparison of HoPE and different RoPE variants in long video understanding. From Table 5.1, we observe that: (1) HoPE consistently outperforms all baselines across nearly all benchmarks, context lengths, and backbone sizes. Specifically, under the 7B model scale and 32k context lengths, HoPE surpasses vanilla RoPE by 2.82, 4.05, and 1.45 on MLVU, LongVideoBench, and Video-MME, respectively. This confirms its effectiveness and generalizability in multimodal long-context modeling. (2) The effectiveness of HoPE scales with backbone size. For instance, when the size of the backbone LLM increases from 2B to 7B, HoPE’s performance gain on LongVideoBench (32k) significantly increases from 0.66 to 4.05 compared to vanilla RoPE. Notably, the performance gap between different methods on the 2B scale is less significant, probably due to the limited capabilities of the backbone LLM. (3) For context lengths under 64k, performance on Video-MME drops substantially, while the impact on MLVU and LongVideoBench is less pronounced. This suggests that extrapolating to extreme context lengths (e.g., up to 8x) remains highly challenging.
Figure 3: Performance comparison on long video retrieval task (V-NIAH). Here, each frame corresponds to 144 tokens. Cell colors indicate model accuracy (red: low, green: high), and the black dotted line marks the training context length (8k).
@@ 5.3 Results on Long Video Retrieval
5.4 Analysis
In this section, we first conduct ablation studies to analyze the effectiveness of each component in HoPE. We then present a comprehensive analysis exploring how different factors, including task type, context length, and the scaling factor of visual tokens, interact and impact model performance.
Ablation Studies. We conduct a series of ablation experiments to evaluate the impact of each component in HoPE and summarize the results in Table 4. According to the results, we observe that :
Figure 4: Ablation results on Video-MME from 8k to 64k. Here, HFA: hybrid frequency allocation, DTS: dynamic temporal scaling.
@@ 5.4 Analysis
+ 3D + HFA 57.66 59.19 59.31 26.98
[…]
(1) The 3D structure effectively improves the performance of vanilla RoPE in multimodal contexts, supporting our Proposition 3.1. (2) Based on the 3D structure, the hybrid frequency allocation (HFA) strategy further enhances long-range semantic modeling, achieving an average improvement of 1.69 across all context lengths. (3) The dynamic temporal scaling (DTS) mechanism facilitates VLMs’ robustness to varying video speeds in real-world scenarios, yielding further performance gain. By combining the above strategies, our HoPE achieves the best overall performance across different context lengths in multimodal long-context modeling.
Table 2: Ablation studies on test-time scaling factor selection. We find that long video understanding generally benefits from larger scaling factors, while long video retrieval yields better results with smaller ones.
Scaling Factor $\mathbf{\gamma}$ LongVideoBench V-NIAH
8k 16k 32k 64k
0.50 54.48 54.29 54.36 52.63 60.89
0.75 54.36 54.97 54.72 52.63 63.56
1.00 54.11 54.48 54.97 52.63 62.67
1.25 54.11 54.84 55.70 52.63 62.67
1.50 54.11 55.09 55.34 51.22 61.78
Impact of Test-time Scaling Factor Selection. We conduct further experiments to investigate how different scaling factors $\gamma$ in our dynamic temporal scaling mechanism impact the performance of video-related tasks. We summarize the results on V-NIAH and LongVideoBench in Table 2. Our main observations are as follows: (1) Long video retrieval generally prefers smaller scaling factors. As shown in Table 2, when we utilize smaller scaling factors $\gamma$ during inference, the performance on V-NIAH improves. We attribute this to the substantial length of 1-hour videos (3,000 frames), which far exceeds the training length (128 frames). In such cases, smaller scaling factors indirectly prevent the spatial position indices from becoming excessively large (see Equation 5), thereby providing a better guarantee for the semantic preference property. Therefore, we set $\gamma=0.75$ for long video retrieval. (2) Long video understanding generally benefits from larger scaling factors. In contrast to retrieval, we find that long video understanding is relatively insensitive to the choice of scaling factor when the input context length is close to the training length. However, as the input length increases, employing larger scaling factors ($\gamma>1$) results in better performance. We hypothesize that while smaller scaling factors help preserve the semantic preference property, larger scaling factors are beneficial for maintaining spatial details (also see Equation 5), which are crucial for complex understanding tasks. This introduces a natural tradeoff between semantic preference and spatial detail preservation. Compared to long video retrieval (3,000 frames, roughly 432k tokens), where extended temporal distances can significantly degrade semantic preference, in long video understanding tasks with context lengths of 16k–32k, the negative impact on semantic preference is relatively small. At the same time, the positive effect of larger scaling factors on capturing spatial details outweighs the semantic preference loss, making larger scaling factors overall more effective for complex video understanding. In our experiments, we set $\gamma=1.5$ for long video understanding.
@@ 6 Related Work
Position Embedding in LLMs. Rotary Position Embedding (RoPE) [23] has become a common choice for position embedding in modern LLMs [24, 25, 26, 44]. As discussed in Section 2, RoPE achieves this success through rotating query and key vectors, encoding relative position information through an absolute positional encoding approach. Despite its success, several works have pointed out that No Position Embedding (NoPE) still works for decoder-only LLMs, arguing that the causal attention mechanism implicitly learns absolute position information [30, 31, 38]. These works even suggest that NoPE outperforms RoPE in out-of-distribution (OOD) scenarios. However, this observation remains unexplored in multimodal settings, where positional encoding strategies may have different implications for cross-modal interactions. Based on Lemma 4.1, we find that incorporating NoPE’s zero frequency strategy indeed improves the length generalization of multimodal RoPE.
Multimodal Position Embedding in VLMs. In VLMs [1, 2, 3, 4, 5], images are first processed by vision encoders and then flattened into 1D tokens. Several early models [1, 4, 5] rely on vanilla RoPE for positional encoding, which distorts spatial-temporal locality (see Section 3) and limits VLMs’ long-context capability. Recently, Qwen2-VL [2] introduced M-RoPE, which extends 1D RoPE to multimodal settings by assigning distinct frequency ranges to different positional components. Specifically, M-RoPE allocates the highest frequencies to the temporal component $t$, while distributing the lower frequencies sequentially to the spatial components $x$ and $y$. Conversely, VideoRoPE [28] allocates the lowest frequencies to $t$ to capture long-range dependencies, achieving stronger length generalization. However, these allocation strategies mainly rely on heuristics, lacking in-depth theoretical analysis. In contrast, our work theoretically analyzes how different frequency allocation strategies impact the performance of multimodal RoPE. By zeroing out low frequencies for temporal modeling, our proposed HoPE provides the strongest theoretical guarantee for long-range semantic modeling. HoPE’s strength is further enhanced by its dynamic temporal scaling of visual tokens, which enables robust temporal learning during training and flexible scaling during inference. By integrating these advantages, HoPE achieves state-of-the-art performance in long video understanding and retrieval tasks, making it well-suited for long context VLMs.
@@ 7 Conclusion
Limitations. While HoPE’s performance gains scale from 2B to 7B backbones, our work does not use larger models or training data. We observe that the performance of all methods degrades significantly at 64k, though HoPE remains the most robust. While these resource-constrained evaluations are essential for uncovering genuine algorithmic benefits of multimodal RoPE, we note that training with more data, particularly long-context data, could further improve length generalization.
@@ A.2 Semantic Preference Property
We now prove that the frequency allocation strategies in current multimodal RoPEs are unreliable in capturing semantic similarities over extended contexts, as defined in Definition 3.1.
Definition 3.1. (Semantic Preference). For any query vector $\mathbf{q}$ and a semantically similar key vector $\mathbf{k}^{\prime}$ that can be expressed as $\mathbf{k}^{\prime}=\mathbf{q}+\delta$ where $\delta$ is a zero-mean perturbation, the attention score with RoPE should satisfy:
@@
2. Monotonicity once the first zero enters. As soon as $\theta\geq\frac{\pi}{2(L-1)}$, the point satisfying $\theta\,\Delta t=\pi/2$ lies in $\{0,\dots,L-1\}$. Each further increase in $\theta$ extends the interval of length $\theta(L-1)$, adding more half-periods of cosine. Each added half-period contains exactly one “negative” region of length $\pi$. Therefore the count $\bigl|\{\Delta:\cos(\theta\Delta)<0\}\bigr|$ (and hence $P_{\neg}(\theta)$) can only stay the same or increase, up to $O(1/L)$ rounding errors on the discrete grid.
3. Limit to one half for large $\theta$. For large $\theta$, the values $\{\theta\,\Delta\}_{\Delta=0}^{L-1}$ become equidistributed mod $2\pi$. Since the negative region $\{\;x\;\bmod\;2\pi\mid\cos x<0\}$ has total length $\pi$ over each $2\pi$-cycle, one finds
@@
Theorem 3.1. Let $X=[x_{1},x_{2},\dots,x_{L}]$ be an input sequence, and let RoPE use any fixed set of temporal frequencies (e.g., highest or lowest). Then there exists a critical length $L_{c}$ such that for all $L\geq L_{c}$, the semantic preference property (Definition 3.1) is violated.
@@
Assume that each component of the query vector $\mathbf{q}$ is independently and identically distributed with mean $\mu$ and variance $\sigma^{2}$. We denote key vector that is semantically similar to $\mathbf{q}$ as $\mathbf{k}^{\prime}=\mathbf{q}+\delta$, where $\delta$ is a zero-mean perturbation. The semantically unrelated key vector $\mathbf{k}$ is independently drawn with the same distribution as $\mathbf{q}$. Let $\Delta t,\Delta x,\Delta y$ denote relative temporal and spatial distances between the query and each key. According to Definition 3.1, the semantic preference property requires that:
@@
where $i_{t},i_{x},i_{y}$ denote dimensions allocated to encode temporal $(t)$, horizontal spatial $(x)$, and vertical spatial $(y)$ information. To satisfy the semantic preference property (Definition 3.1), the expected attention between a query and its semantically similar key should remain higher than that for an unrelated key, regardless of their relative distance. This implies the following condition must hold universally:
@@
However, despite using the lowest frequencies, VideoRoPE still fails to guarantee that the semantic preference property holds for all context lengths (Equation 10). Let VideoRoPE allocate only the smallest frequency to the temporal dimensions, instead of $|i_{t}|$ smallest frequencies:
@@
$2\sigma^{2}\ |i_{t}|\cos(\Delta t\ \cdot\theta_{\min})<0.$
This single counterexample $(\Delta t,\Delta x,\Delta y)$ violates the semantic preference condition, since no further temporal frequencies are available to “rescue” the sum. Therefore, despite using the lowest frequencies for temporal modeling, VideoRoPE still fails to guarantee the semantic preference property. In conclusion, all frequency allocation strategies in current multimodal RoPEs fail to maintain the semantic preference property in Definition 3.1, completing the proof. ∎